# RAZOR: Sharpening Knowledge by Cutting Bias with Unsupervised Text Rewriting

### v1

@@ 1 Introduction
With the progression of instructional learning, prominent LLMs such as ChatGPT (Ouyang et al. 2024), ChatGLM (GLM et al. 2024), and Vicuna (Zheng et al. 2023) are extensively utilized across diverse domains owing to their versatility and outstanding performance (Chang et al. 2024).
However, their utility is significantly constrained by the high training and API licensing costs. As a result, numerous researchers in domains like healthcare, education, and e-commerce adhere to the conventional approach of using open-source language models followed by fine-tuning (Strubell, Ganesh, and McCallum 2019).
LLMs absorb linguistic features from extensive corpora during pre-training by predicting specific lexical units. This foundational learning phase is then followed by fine-tuning domain-specific data, which enhances their ability to effectively perform domain-specific tasks.
@@
However, unlike linguistic features, surface biases are reported as commonly existing in manually curated datasets and can detrimentally affect the fine-tuning phase (Jiménez, Mendes Oliveira, and Soares 2020). In other words, models may achieve untruthfully elevated performance during training by exploiting specific shortcuts. For instance, analyses of the FEVER dataset (Thorne et al. 2018), used for fact verification tasks, revealed a pronounced association between the occurrence of negations within the claim and the corresponding label “refutes” (Karimi Mahabadi, Belinkov, and Henderson 2020). This suggests that models might erroneously generalize that claims containing negations are invariably fake.
@@
The main contributions of this paper are as follows:
@@
  • To the best of our knowledge, we are the first to formalize the definition of shortcuts in relation to an underlying classifier and the ground truth of the sentences. Our theoretical findings help us understand RAZOR’s effectiveness in mitigating spurious correlations in real-world datasets without prior knowledge.
@@ 2 Related Work
Wu et al. (2022) present a data generation approach aimed at mitigating spurious correlations in natural language inference datasets through diverse data generation. The authors fine-tune GPT-2 as their data generation model to sequentially engender an instance’s premise, label, and hypothesis. Additionally, they apply unlikelihood training to mitigate label inconsistency phenomena, improving the label quality at generation time. Finally, they add a consistency step (Bartolo et al. 2021; Lewis et al. 2021) – i.e., using the model’s confidence for filtering purposes – to improve the quality of the generated dataset further. After this process, the authors reject samples that contribute to the high spurious correlations between task-independent features of the samples and their labels via z-filtering and produce a debiased dataset. Bras et al. (2020) propose an iterative greedy algorithm that adversarially filters out spurious biases from data by relying on AFLite (Sakaguchi et al. 2020). Interestingly, the authors aim to mitigate performance leakage when datasets present spurious biases. [^2] Spurious biases in datasets make benchmarks artificially easier as models learn to overly rely on these biases instead of learning to generalize. This also hurts out-of-domain generalization. CrossAug (Lee et al. 2021) reduces dataset bias by creating contrastive pairs in fact-checking tasks. It uses a two-stage pipeline where the first stage generates negative claims by altering positive claims, and the second stage modifies the evidence to support these negative claims. In this way, CrossAug enhances a classification model’s ability to reason more accurately based on the provided evidence piece of text. Finally, CLEVER (Xu et al. 2023a) leverages counterfactual inference for debiasing in fact-checking tasks. Instead of relying on data augmentation or training-phase adjustments, CLEVER focuses on the inference stage. It does so by training two models, a claim-evidence fusion model and a claim-only model, and then subtracting the output of the claim-only model from the fusion model to isolate and remove biases. EDA (Wei and Zou 2019) involves four simple techniques: synonym replacement, random insertion, random swap, and random deletion, all modifying the text while preserving its original meaning. These operations generate diverse training examples, helping models generalize better by exposing them to varied linguistic patterns. The approach is designed to be easy to implement and computationally inexpensive. EDA mitigates overfitting and improves model robustness by focusing on random but controlled transformations. Schuster et al. (2019) demonstrate that models trained solely on claims (without evidence) can achieve high accuracy due to dataset biases, particularly the presence of giveaway phrases that correlate with specific labels. To mitigate this issue, they create a “Symmetric Test Set”, which eliminates these biases by manually generating balanced claim-evidence pairs. Additionally, they propose a regularization method – namely ReW – that reweights training examples to reduce the influence of these biases during model training. Mahabadi, Belinkov, and Henderson (2020) propose “Product of Experts” – namely PoE – which combines predictions from a bias-only model and the main model, focusing the training on examples that are not biased.
@@
Lately, augmenting models with counterfactual data has become a favored method for reducing spurious correlations and enhancing model robustness. Kaushik and Lipton (2018) initially employed human workers to generate counterfactual examples for augmentation. Their findings indicate that counterfactually augmented data effectively mitigates spurious patterns in training data. However, this approach is costly, time-intensive, and often results in simple perturbations. Wu et al. (2021) and Ross et al. (2022) leverage text generation models to produce counterfactual data. These models require fine-tuning with specific perturbation types and have various limitations, among which the untargeted and unlabelled generation process and the perturbation types being restricted. To introduce new perturbation types, the models need retraining. Unlike the previous two methods, Chen et al. (2023) develop DISCO, a framework for distilling counterfactuals with large language models to help distinguish between spurious and genuine correlations. DISCO engineers prompts to generate phrasal perturbations with a large general language model. Then, a task-specific teacher model filters these generations to distill high-quality counterfactual data to enrich the original training set. Lastly, Xu et al. (2023b) focus on counterfactual debiasing for fact verification, generating counterfactual scenarios to identify and mitigate spurious correlations in verification tasks.
@@ 3.1 Problem Formulation
We consider binary classifications of short text documents – e.g., sentiment classification on reviews. Let us assume that we have a set of i.i.d. labelled samples – i.e., dataset – $\mathcal{D}=\{d_{1},\dots,d_{n}\}$. Each sample in $\mathcal{D}$ is a document, i.e., a sequence of tokens, $d_{i}=\{t_{1},t_{2},\dots,t_{m}\}$ with a corresponding label $y_{i}\in\mathcal{Y}$ where w.l.o.g. $\mathcal{Y}=\{0,1\}$. To classify a document $d_{i}$, we first transform it into a feature vector $\mathbf{x}_{i}$ via a designated transformation function $g:\mathcal{D}\rightarrow\mathcal{X}\subseteq\mathbb{R}^{u}$, where $\mathcal{X}$ is the domain of features. Then, $\mathbf{x}_{i}$ is classified by a classification function $f_{\theta}:\mathcal{X}\rightarrow\mathcal{Y}$ with model parameters $\theta$. The parameters $\theta$ are typically estimated from $\mathcal{D}$ by optmizing a certain loss function $\theta^{*}\leftarrow\arg\min_{\theta}\mathcal{L}(\mathcal{D},\theta)$.
Generally, the classification function $f_{\theta}$ adopted in NLP is a deep learning model (see among others (Naseem et al. 2020; Li et al. 2024a)). Alas, as noticed in (Izmailov et al. 2022), deep classifiers are known to rely on spurious features – i.e., patterns that are correlated with the target on the training data but not inherently relevant to the learning problem, such as the image backgrounds when classifying the foreground. In NLP problems, such as sentiment classification tasks, certain tokens (e.g., words) can be associated with positive/negative classes due to their co-occurrence with the labels in the training data. Inspired by (Wang et al. 2022), in Fig. 1, we show how the classifier exploits the shortcut terms Spielberg and New York Subway to misclassify the sentences.
While shortcut mitigation has been extensively studied, to the best of our knowledge, there is a lack of a uniform definition of what a shortcut represents. Hence, we provide the reader with a formal definition of shortcut tokens.
Notice that each document $d_{i}$, before being transformed into a feature vector $x_{i}$, gets tokenized into a sequence of tokens $\{t_{1},t_{2},\dots,t_{m}\}$, e.g., at a word-level, character-level, or sub-word level. Lastly, assume that $\Phi:\mathcal{X}\rightarrow\mathcal{Y}$ represents the ground truth function.
@@
Attention mechanism. For completeness, let us remind the reader how a single-head attention mechanism works in a transformer. Given a document $d_{i}=\{t_{1},t_{2},\dots,t_{m}\}$, the attention mechanism $h:\mathcal{X}\rightarrow\mathbb{R}^{\ell}$ typically involves the following steps:
  • Linear transformation – Each token embedding $g(t_{j})$ is linearly transformed into query $\mathbf{Q}_{j}\in\mathbb{R}^{\ell}$, key $\mathbf{K}_{j}\in\mathbb{R}^{\ell}$, and value $\mathbf{V}_{j}\in\mathbb{R}^{\ell}$ vectors.
  • Attention scores calculation – The attention score $\alpha_{jk}$ between tokens $t_{j}$ and $t_{k}$ is computed using the scaled [^3] In the original formulation, the dot-product between the query and key vectors is scaled by $\sqrt{\ell}$ to maintain gradient stability. dot-product $\alpha_{jk}=\mathbf{Q}_{j}\cdot\mathbf{K}_{k}^{\top}$.
  • Softmax operation – The attention scores are passed through a softmax function to obtain the attention weights
    $\alpha^{\prime}_{jk}=\frac{\text{exp}(\alpha_{jk})}{\sum_{j^{\prime}=1}^{\ell}\text{exp}(\alpha_{jj^{\prime}})}$
  • Weighted sum – The output for token $t_{j}$ is a weighted sum of the value vectors $\mathbf{V}_{k}$ using the attention weights
    $\mathbf{z}_{j}=\sum_{k=1}^{\ell}\alpha^{\prime}_{jk}\mathbf{V}_{k}$
In this work, we assume that $h(g(t_{j}))=\mathbf{z}_{j}\;\forall t_{j}\in d_{i}$.
Essential vs. non-essential tokens. According to Def. 1, we have two sets of tokens for each document $d_{i}$: i.e., the set of “essential” tokens $\hat{d}_{i}$, and that of “non-essential” ones $d_{i}\backslash\hat{d}_{i}$. The essential tokens are the ones that contribute the most (i.e., their attention scores are high) to the prediction. In contrast, non-essential tokens do not significantly contribute to the prediction (Vaswani et al. 2017). Because the prediction does not change if we use the entire set of tokens in the document $d_{i}$, intuitively, one can conclude that the attention scores of the set of non-essential are lower than those of the essential set of tokens (see Lemma 1).
@@
We invite the reader to check Appendix A for the proof of Lemma 1. In practice, we observe that shortcuts are much smaller than the rest of the tokens in a sentence, which paves the way for efficient methods to identify the subset of tokens that represent shortcuts
Figure 2: An example of RAZOR’s application in the fact-checking task. The task aims to determine whether a piece of evidence from Wikipedia “supports” () or “refutes” ($\times$) a claim. The instance here is sampled from the FEVER dataset, where a negation word ”not” has been reported to exhibit a spurious correlation with the $\times$.
@@ 3.2 Shortcut Identification
To find $\widehat{\mathcal{D}}$, we define a surface feature extraction function $\hat{g}:\mathcal{D}\rightarrow\mathbb{R}^{\lambda}$ which maps all samples into a shortcut space. Notably, the specific types of shortcuts are strongly task-dependent. In this work, without any prior knowledge, we provide a general identification pattern with the two essential factors by following Def. 1:
  1. 1.
    The significance of tokens.
  2. 2.
    The (relative) positions of tokens.
@@
In other words, given a document $d_{i}\in\mathcal{D}$, and a ground truth function $\Phi$, as introduced in Def. 1, $\varphi(d_{i},\Phi)$ produces a set of documents whose label is different with $\Phi(d_{i})$. In this way, we define a shortcut score $\gamma:\mathcal{D}\rightarrow\mathbb{R}$ as in Eq. 12.
@@ 3.3 Rewriting and Filtering
Once, the rewriting process is completed for a particular $d_{i}$, we obtain a set of generated documents $\varphi(d_{i})=\{d_{i}^{*}\;|\;d_{i}^{*}=G_{\alpha}(d_{i},\Phi(d_{i}))\}$. From $\varphi(d_{i})$, we choose the one scoring the least according to Eq. 12 and insert it in the original dataset $\mathcal{D}$. The original document $d_{i}$ is removed from the “rewritten” dataset as we aim to mitigate the shortcuts that it might have contained. In this way, RAZOR maintains the same size as the original dataset throughout the rewriting process. We repeat this process until the Kullback-Leibler divergence between the embeddings in the surface feature space of the documents of each label is minimal (see Eq. 14).
$\begin{gathered}\text{minimize}\sum_{d_{i}\in\mathcal{D}_{0}}\sum_{d_{j}\in\mathcal{D}_{1}}\text{KL}(\hat{g}(d_{i}),\hat{g}(d_{j}))\\ \text{s.t. }\mathcal{D}_{y}=\{d\;|\;d\in\mathcal{D}\;\land\;\Phi(d)=y\}\end{gathered}$ (14)
@@ 4 Experiments
Table 1: Comparison of RAZOR with SoTA methods on the FEVER and FEVER-Adversarial datasets – the datasets have the same training set, but different test sets. We show the performances of BERT and RoBERTa classifiers in terms of Accuracy and F1 scores after the shortcuts within the dataset have been modified from each debiasing method, including RAZOR (ours). Original means that the original training set has not been modified, and we report the performances of the two classifiers as is.
@@
RAZOR(with Llama) 87.97$\pm$1.20 87.22$\pm$2.03 90.59$\pm$2.32 89.11$\pm$2.56 66.03$\pm$1.95 69.61$\pm$1.88 68.77$\pm$3.01 73.76$\pm$2.50
RAZOR (with ChatGPT) 88.73$\pm$2.23 88.39$\pm$2.20 90.45$\pm$1.99 90.22$\pm$2.11 66.66$\pm$1.98 73.78$\pm$1.37 69.21$\pm$1.64 76.08$\pm$2.02
Table 2: RAZOR’s impact on the BERT and RoBERTa classifiers on the NLI datasets in terms of Accuracy and F1 score. Note that we assess the performances of BERT and RoBERTa across different train-test set combinations – e.g., training on MNLI and testing on SNLI – to mitigate that the original test set contains similar shortcuts as the original training set. Within parentheses, we indicate the training set; outside of them, we indicate the test set.
@@
w/ RAZOR 88.50$\pm$1.32 87.90$\pm$1.80 84.75$\pm$1.72 85.07$\pm$2.44 74.10$\pm$3.20 67.86$\pm$3.57 93.20$\pm$1.11 93.16$\pm$1.26
$\Delta$ +2.08 +1.44 +2.04 +0.78 +5.63 +8.14 +1.69 +1.55
@@
w/ RAZOR 91.15$\pm$1.26 91.14$\pm$1.48 91.45$\pm$1.29 91.42$\pm$1.68 86.50$\pm$2.20 86.22$\pm$1.72 95.15$\pm$0.70 95.17$\pm$0.98
$\Delta$ +0.33 +0.42 +1.27 +0.37 +1.34 +1.52 +0.11 +0.17
Table 3: RAZOR with Llama
MNLI (MNLI) SNLI (MNLI) MNLI (SNLI) SNLI (SNLI)
Accuracy F1 Accuracy F1 Accuracy F1 Accuracy F1
w/ RAZOR 86.80$\pm$1.32 86.88$\pm$1.80 84.10$\pm$1.52 83.51$\pm$2.00 74.25$\pm$2.92 68.73$\pm$3.14 92.90$\pm$1.27 92.81$\pm$1.32
$\Delta$ +0.12 +0.04 +1.26 -1.06 +5.85 +9.56 +1.36 +1.17
w/ RAZOR 91.55$\pm$1.38 91.45$\pm$1.70 90.90$\pm$1.29 91.13$\pm$1.68 85.50$\pm$1.98 85.76$\pm$2.24 95.60$\pm$0.70 95.71$\pm$0.98
$\Delta$ +0.77 +0.77 +0.66 +0.05 +0.18 +0.98 +0.58 +0.74

### v2

@@ 1 Introduction
With the progression of instructional learning, prominent LLMs such as ChatGPT (Ouyang et al. 2024) and Vicuna (Zheng et al. 2023) are extensively utilized across diverse domains owing to their versatility and outstanding performance (Chang et al. 2024). However, their utility is significantly constrained by the high training and API licensing costs. As a result, numerous researchers in domains like healthcare, education, and e-commerce adhere to the conventional approach of using open-source language models followed by fine-tuning (Strubell, Ganesh, and McCallum 2019).
@@
LLMs absorb linguistic features from extensive corpora during pre-training by predicting specific lexical units. This foundational learning phase is then followed by fine-tuning domain-specific data, enhancing their ability to perform domain-specific tasks effectively.
Unlike linguistic features, surface biases are reported as commonly existing in manually curated datasets and can detrimentally affect the fine-tuning phase (Jiménez, Mendes Oliveira, and Soares 2020). In other words, models may achieve untruthfully elevated performance during training by exploiting specific shortcuts. For instance, analyses of the FEVER dataset (Thorne et al. 2018), used for fact verification tasks, revealed a pronounced association between the occurrence of negations within the claim and the corresponding label “refutes” (Karimi Mahabadi, Belinkov, and Henderson 2020). This suggests that models might erroneously generalize that claims containing negations are invariably fake.
@@
The main contributions of this paper are as follows: [^2] Our code is available at https://github.com/ShuoYangtum/RAZOR. The appendix sections can be found at https://arxiv.org/abs/2412.07675.
@@
  • To the best of our knowledge, we are the first to formalize the definition of shortcuts about an underlying classifier and the ground truth of the sentences. Our theoretical findings help us understand RAZOR’s effectiveness in mitigating spurious correlations in real-world datasets without prior knowledge.
@@ 2 Related Work
Wu et al. (2022) present a data generation approach aimed at mitigating spurious correlations in natural language inference datasets through diverse data generation. The authors fine-tune GPT-2 as their data generation model to sequentially engender an instance’s premise, label, and hypothesis. Additionally, they apply unlikelihood training to mitigate label inconsistency phenomena, improving the label quality at generation time. Finally, they add a consistency step (Bartolo et al. 2021; Lewis et al. 2021) – i.e., using the model’s confidence for filtering purposes – to improve the quality of the generated dataset further. After this process, the authors reject samples that contribute to the high spurious correlations between task-independent features of the samples and their labels via z-filtering and produce a debiased dataset. Bras et al. (2020) propose an iterative greedy algorithm that adversarially filters out spurious biases from data by relying on AFLite (Sakaguchi et al. 2020). Interestingly, the authors aim to mitigate performance leakage when datasets present spurious biases. [^3] Spurious biases in datasets make benchmarks artificially easier as models learn to overly rely on these biases instead of learning to generalize. This also hurts out-of-domain generalization. CrossAug (Lee et al. 2021) reduces dataset bias by creating contrastive pairs in fact-checking tasks. It uses a two-stage pipeline where the first stage generates negative claims by altering positive claims, and the second stage modifies the evidence to support these negative claims. In this way, CrossAug enhances a classification model’s ability to reason more accurately based on the provided evidence piece of text. Finally, CLEVER (Xu et al. 2023a) leverages counterfactual inference for debiasing in fact-checking tasks. Instead of relying on data augmentation or training-phase adjustments, CLEVER focuses on the inference stage. It does so by training two models, a claim-evidence fusion model, and a claim-only model, and then subtracting the output of the claim-only model from the fusion model to isolate and remove biases. EDA (Wei and Zou 2019) involves four simple techniques: synonym replacement, random insertion, random swap, and random deletion, all modifying the text while preserving its original meaning. These operations generate diverse training examples, helping models generalize better by exposing them to varied linguistic patterns. The approach is designed to be easy to implement and computationally inexpensive. EDA mitigates overfitting and improves model robustness by focusing on random, controlled transformations. Schuster et al. (2019) demonstrate that models trained solely on claims (without evidence) can achieve high accuracy due to dataset biases, particularly the presence of giveaway phrases that correlate with specific labels. To mitigate this issue, they create a “Symmetric Test Set”, which eliminates these biases by manually generating balanced claim-evidence pairs. Additionally, they propose a regularization method – namely ReW – that reweights training examples to reduce the influence of these biases during model training. Mahabadi, Belinkov, and Henderson (2020) propose “Product of Experts” – namely PoE – which combines predictions from a bias-only model and the main model, focusing the training on examples that are not biased.
@@
Lately, augmenting models with counterfactual data has become a favored method for reducing spurious correlations and enhancing model robustness. Kaushik and Lipton (2018) initially employed human workers to generate counterfactual examples for augmentation. Their findings indicate that counterfactually augmented data effectively mitigates spurious patterns in training data. However, this approach is costly, time-intensive, and often results in simple perturbations. Wu et al. (2021) and Ross et al. (2022) leverage text generation models to produce counterfactual data. These models require fine-tuning with specific perturbation types and have various limitations, including the untargeted and unlabelled generation process and the restricted perturbation types. To introduce new perturbation types, the models need retraining. Unlike the previous two methods, Chen et al. (2023) develop DISCO, a framework for distilling counterfactuals with large language models to help distinguish between spurious and genuine correlations. DISCO does prompt engineering to generate phrasal perturbations with a large general language model. Then, a task-specific teacher model filters these generations to distill high-quality counterfactual data to enrich the original training set. Lastly, Xu et al. (2023b) focus on counterfactual debiasing for fact verification, generating counterfactual scenarios to identify and mitigate spurious correlations in verification tasks.
@@ 3.1 Problem Formulation
We consider binary classifications of short text documents, e.g., sentiment classification of reviews. Let us assume that we have a set of i.i.d. labelled samples – i.e., dataset – $\mathcal{D}=\{d_{1},\dots,d_{n}\}$. Each sample in $\mathcal{D}$ is a document, i.e., a sequence of tokens, $d_{i}=\{t_{1},t_{2},\dots,t_{m}\}$ with a corresponding label $y_{i}\in\mathcal{Y}$ where w.l.o.g. $\mathcal{Y}=\{0,1\}$. To classify a document $d_{i}$, we first transform it into a feature vector $\mathbf{x}_{i}$ via a designated transformation function $g:\mathcal{D}\rightarrow\mathcal{X}\subseteq\mathbb{R}^{u}$, where $\mathcal{X}$ is the domain of features. Then, $\mathbf{x}_{i}$ is classified by a classification function $f_{\theta}:\mathcal{X}\rightarrow\mathcal{Y}$ with model parameters $\theta$. The parameters $\theta$ are typically estimated from $\mathcal{D}$ by optmizing a certain loss function $\theta^{*}\leftarrow\arg\min_{\theta}\mathcal{L}(\mathcal{D},\theta)$.
Generally, the classification function $f_{\theta}$ adopted in NLP is a deep learning model (see among others (Naseem et al. 2020; Li et al. 2024a)). Alas, as noticed in (Izmailov et al. 2022), deep classifiers are known to rely on spurious features – i.e., patterns that are correlated with the target on the training data but not inherently relevant to the learning problem, such as the image backgrounds when classifying the foreground. In NLP problems, such as sentiment classification tasks, certain tokens (e.g., words) can be associated with positive/negative classes due to their co-occurrence with the labels in the training data. Inspired by (Wang et al. 2022), in Fig. 1, we show how the classifier exploits the shortcut terms Spielberg and New York Subway to misclassify the sentences. While shortcut mitigation has been extensively studied, to the best of our knowledge, there is a lack of a uniform definition of what a shortcut represents. Hence, we provide the reader with a formal definition of shortcut tokens. Notice that each document $d_{i}$ gets tokenized into a sequence of tokens $\{t_{1},t_{2},\dots,t_{m}\}$, e.g., at a word-level, character-level, or sub-word level before being transformed into a feature vector $x_{i}$. Lastly, assume that $\Phi:\mathcal{X}\rightarrow\mathcal{Y}$ represents the ground truth function.
@@
Attention mechanism. Let $h:\mathcal{X}\rightarrow\mathbb{R}^{\ell}$ be the attention mechanism that, given in input a document $d_{i}=\{t_{1},t_{2},\dots,t_{m}\}$, produces attention weights $\mathbf{z}_{j}$ for each token $t_{j}\in d_{i}$. In this work, we assume that $h(g(t_{j}))=\mathbf{z}_{j}\;\forall t_{j}\in d_{i}$.
Figure 2: An example of RAZOR’s application in the fact-checking task. The task aims to determine whether a piece of evidence from Wikipedia supports or refutes a claim. The instance here is sampled from the FEVER dataset, where a negation word ”not” has been reported to exhibit a spurious correlation with the class refutes.
Essential vs. non-essential tokens. According to Definition 1, we have two sets of tokens for each document $d_{i}$: i.e., the set of “essential” tokens $\hat{d}_{i}$, and that of “non-essential” ones $d_{i}\backslash\hat{d}_{i}$. The essential tokens are the ones that contribute the most (i.e., their attention scores are high) to the prediction. In contrast, non-essential tokens do not significantly contribute to the prediction (Vaswani et al. 2017). Because the prediction does not change if we use the entire set of tokens in the document $d_{i}$, intuitively, one can conclude that the attention scores of the set of non-essential are lower than those of the essential set of tokens (see Lemma 1).
@@
We invite the reader to check Appendix A for the proof of Lemma 1. In practice, we observe that shortcuts are much smaller than the rest of the tokens in a sentence, which paves the way for efficient methods to identify the subset of tokens that represent shortcuts.
@@ 3.2 Shortcut Identification
To find $\widehat{\mathcal{D}}$, we define a surface feature extraction function $\hat{g}:\mathcal{D}\rightarrow\mathbb{R}^{\lambda}$ which maps all samples into a shortcut space. Notably, the specific types of shortcuts are strongly task-dependent. In this work, without any prior knowledge, we provide a general identification pattern with the two essential factors by following Definition 1: i.e., (1) the significance of tokens, and (2) the (relative) positions of tokens.
@@
In other words, given a document $d_{i}\in\mathcal{D}$, and a ground truth function $\Phi$, as introduced in Definition 1, $\vartheta(d_{i},\Phi)$ produces a set of documents whose label is different with $\Phi(d_{i})$. In this way, we define a shortcut score $\gamma:\mathcal{D}\rightarrow\mathbb{R}$ as in Eq. 12.
@@ 3.3 Rewriting and Filtering
Once, the rewriting process is completed for a particular $d_{i}$, we obtain a set of generated documents $\varphi(d_{i})=\{d_{i}^{*}\;|\;d_{i}^{*}=G_{\alpha}(d_{i},\Phi(d_{i}))\}$. From $\varphi(d_{i})$, we choose the one scoring the least according to Eq. 12 and insert it in the original dataset $\mathcal{D}$. The original document $d_{i}$ is removed from the “rewritten” dataset as we aim to mitigate the shortcuts that it might have contained. In this way, RAZOR maintains the same size as the original dataset throughout the rewriting process. We repeat this process until the cosine similarity between the embeddings in the surface feature space of the documents of each label is maximal (see Eq. 14).
$\begin{gathered}\text{maximize}\sum_{d_{i}\in\mathcal{D}_{0}}\sum_{d_{j}\in\mathcal{D}_{1}}\text{cos}(\hat{g}(d_{i}),\hat{g}(d_{j}))\\ \text{s.t. }\mathcal{D}_{y}=\{d\;|\;d\in\mathcal{D}\;\land\;\Phi(d)=y\}\end{gathered}$ (14)
@@ 4 Experiments
Table 1: Comparison of RAZOR with SoTA methods on the FEVER and FEVER-Adversarial datasets – the datasets have the same training set but different test sets. We show the performance of the classifiers in terms of accuracy, and F1 scores after the shortcuts within the dataset have been modified using each debiasing method, including RAZOR (ours). Original means that the original training set has not been modified, and we report the performances of the two classifiers as is.
@@
RAZOR (w/ LLaMA) 87.97$\pm$1.20 87.22$\pm$2.03 90.59$\pm$2.32 89.11$\pm$2.56 66.03$\pm$1.95 69.61$\pm$1.88 68.77$\pm$3.01 73.76$\pm$2.50
RAZOR (w/ GPT) 88.73$\pm$2.23 88.39$\pm$2.20 90.45$\pm$1.99 90.22$\pm$2.11 66.66$\pm$1.98 73.78$\pm$1.37 69.21$\pm$1.64 76.08$\pm$2.02
Table 2: RAZOR’s impact on the classifiers on the NLI datasets in terms of Accuracy and F1 score. Note that we assess the performances of the classifiers across different train-test set combinations – e.g., training on MNLI and testing on SNLI – to mitigate the fact that the original test set contains similar shortcuts as the original training set. Within parentheses, we indicate the training set; outside of them, we indicate the test set.
@@
w/ RAZOR (GPT) 88.50$\pm$1.32 87.90$\pm$1.80 84.75$\pm$1.72 85.07$\pm$2.44 74.10$\pm$3.20 67.86$\pm$3.57 93.20$\pm$1.11 93.16$\pm$1.26
w/ RAZOR (LLaMA) 86.80$\pm$1.32 86.88$\pm$1.80 84.10$\pm$1.52 83.51$\pm$2.00 74.25$\pm$2.92 68.73$\pm$3.14 92.90$\pm$1.27 92.81$\pm$1.32
$\Delta$ (GPT) +2.08 +1.44 +2.04 +0.78 +5.63 +8.14 +1.69 +1.55
$\Delta$ (LLaMA) +0.12 +0.04 +1.26 -1.06 +5.85 +9.56 +1.36 +1.17
@@
w/ RAZOR (GPT) 91.15$\pm$1.26 91.14$\pm$1.48 91.45$\pm$1.29 91.42$\pm$1.68 86.50$\pm$2.20 86.22$\pm$1.72 95.15$\pm$0.70 95.17$\pm$0.98
w/ RAZOR (LLaMA) 91.55$\pm$1.38 91.45$\pm$1.70 90.90$\pm$1.29 91.13$\pm$1.68 85.50$\pm$1.98 85.76$\pm$2.24 95.60$\pm$0.70 95.71$\pm$0.98
$\Delta$ (GPT) +0.33 +0.42 +1.27 +0.37 +1.34 +1.52 +0.11 +0.17
$\Delta$ (LLaMA) +0.77 +0.77 +0.66 +0.05 +0.18 +0.98 +0.58 +0.74
DistilBERT 83.50$\pm$1.23 82.97$\pm$1.77 77.23$\pm$1.85 77.38$\pm$1.86 65.55$\pm$2.30 69.90$\pm$1.78 88.82$\pm$1.22 87.78$\pm$1.67
w/ RAZOR (GPT) 85.14$\pm$1.66 85.20$\pm$1.84 80.22$\pm$1.90 78.82$\pm$2.21 68.58$\pm$2.16 72.80$\pm$1.83 91.82$\pm$0.88 91.80$\pm$1.03
w/ RAZOR (LLaMA) 84.00$\pm$1.35 83.98$\pm$1.78 77.35$\pm$1.43 77.62$\pm$1.99 66.33$\pm$2.00 71.30$\pm$1.85 90.30$\pm$1.26 90.23$\pm$1.97
$\Delta$ (GPT) +1.96 +2.69 +3.87 +1.86 +4.62 +4.14 +3.38 +4.58
$\Delta$ (LLaMA) +0.60 +1.22 +0.16 +0.31 +1.19 +2.00 +1.67 +2.79